The price of one-year, two-year, and three-year zero-coupon Treasury bond are currently $980, $860, and $780 respectively. The government plans to issue a new three-year annual coupon bond with a coupon rate of 4%. The face value of all the bonds are $1,000. What is the fair price of the new coupon bond? (show work)

A. $1000

B. $884.80

C. $780

D. $1,040

E. $860
Answer #1

Zero coupon bond value = M/(1+i)^n

Where

M= maturity value or face value = 1000

i = required interest rate = 4%=0.04

n= years until maturity = 3

Zero coupon bond value = 1000/(1.04)^3

= 1000/1.124864

= 888.996 (approximately 884.8)

Answer : B